OpenStudy (anonymous):

What is the solution to the system of equations? x + y=5 x-2y=2 ( , )

OpenStudy (anonymous):

@Godlovesme plz help

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@mathmath333 plz help

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

@Sadworld plz help

Directrix (directrix):

Do you have a candidate for an answer @TeenWolfGirl

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ x + y=5----\color{red}{(1)}\\~\\ x-2y=2------\color{red}{(2)}\\~\\ \color{red}{1-2}\\~\\ x+y-(x-2y)=5-2\\~\\ x+y-x+2y=3\\~\\ 3y=3\\~\\ y=1 }\end{align}\) now substitute the value of \(y\) and get \(x\)

OpenStudy (anonymous):

You can use elimination or substitution to solve the system of equations

Directrix (directrix):

x + y=5 ---> x + y = 5 x- 2y =2 ---> -x +2y = -2 Multiplied 2nd equation by -1 and will add to eliminate ======= =========

Directrix (directrix):

@TeenWolfGirl What is the ordered pair?

OpenStudy (anonymous):

-1,1

Directrix (directrix):

No, not quite. Your x coordinate is wrong. Try again.

OpenStudy (anonymous):

sorry not good at this....ummm.....

OpenStudy (anonymous):

4,1 i believe

Directrix (directrix):

That is what I got.

OpenStudy (anonymous):

:)

Directrix (directrix):

Don't forget to encase your answers in this ordered pair notation. ( , )

OpenStudy (anonymous):

ok :) sorry forgot